Good day! Can you please help me with this problem? Thank you!
In triangle ABC, Angle B is 2x-3 and Angle C is 3x+1, what is the measure of Angle A?

Answer by shweta(56)

In triangle ABC, sum of all the angles is 180 degree
Angle A+ Angle B+ Angle C =180
Angle A+ 2x-3+3x+1 =180
Angle A+ 5x-2=180
Angle A=180-5x+2
Angle A=182-5x